A 5.0 kg bowling ball and a 0.15 kg lacrosse ball are 50 cm apart. What gravitation force do they exert on each other?
1
Expert's answer
2018-04-17T11:28:12-0400
F=G*m1*m2/r^2 = 6.67×10^(−11) * 5 * 0.15 / (0.5^2) = 2*10^(-10) N